OpenStudy (waheguru):

How do you solve this 4x^2=8x-1 Finding the roots

OpenStudy (aaronq):

move everything to once side and factor it (or use the quadratic formula).

OpenStudy (waheguru):

I did that but my answer does not match the textbook

OpenStudy (aaronq):

\(4x^2=8x-1\) \(4x^2-8x+1=0\) use the quadratic formula

OpenStudy (waheguru):

Hmm, Why cannt it be -4x^2... you have moves the other digits over is there a reason

OpenStudy (aaronq):

it can be \(4x^2−8x+1=0\) or \(-4x^2+8x-1=0\) you need to plug into the quadratic formula now

OpenStudy (aaronq):

close this please